25 June 2000  Page 1 of 1  J3/00-236

Subject: Issue 273, and other work on passed-object dummy arguments From: Van Snyder

1 Introduction

At meeting 153, /data thought a constraint was an inappropriate place for a definition, so we moved the definition of passed-object dummy arguments from the constraints in 4.5.1 (at [43:7-10, 43:25-28 and 43:38-42]) to a new section 4.5.1.6. We overlooked that there remained two constraints on page 43 that referred to the passed-object dummy argument, for which the definition is now several pages distant, and forward. Issue 273 finds the newly-introduced wording in 4.5.1.6 confusing.

2 Edits

Edits refer to 00-007r2. Page and line numbers are displayed in the margin. Absent other instructions, a page and line number or line number range implies all of the indicated text is to be replaced by immediately following text, while a page and line number followed by + indicates that immediately following text is to be inserted after the indicated line. Remarks for the editor are noted in the margin, or appear between [ and ] in the text.

[Editor: After “in” add “4.5.1.6 and”.] 43:11

[Editor: Delete (moved to [48:35-38]).] 43:12-13

[Editor: After “in” add “4.5.1.6 and”.] 43:29

[Editor: After “in” add “4.5.1.6 and”.] 43:43

[Editor: Delete (moved to [48:35-38]).] 43:44-45

[Editor: Delete “same”. Replace “as the type name” by “named by the type-name”.] 48:33

[Editor: Replace issue 273 by:] 48:35-38 The following constraint applies to syntax rules R435, R436, R441 and R442. Constraint: All of the nonkind type parameters of the passed-object dummy argument shall be

assumed.